OpenStudy (anonymous):

-(4)/((w+1)(w+3))=-3+(2)/(w+3)

OpenStudy (anonymous):

\[\frac{-1}{(w+1)(w+3)}=-3+\frac{2}{(w+3)}\]\[\frac{-1}{(w+1)}=-3(w+3)+2\]\[\frac{-1}{(w+1)}=-3w-9+2\]\[\frac{-1}{(w+1)}=(-3w-7)\]\[-1=(w+1)(-3w-7)\]\[-1=w^2-7w-3w-7\]\[0=-3w^2-10w-6\] Use the quadratic rule to solve for w. \[w = \frac{1}{3} (-5-\sqrt(7))\]\[w = \frac{1}{3} (\sqrt(7)-5)\]